Q1.Sections 1 and 2 are at the beginning and end of the bend of the 200 mm diameter pipe in which the quantity of flow is 0.28 m3/s. The angle of deflection of the water is 40°. Calculate the force that the liquid exerts on the bend if the pressure in the pipe is 50 kPa. Assume no loss of pressure round the bend

Please help me
1
Expert's answer
2019-04-10T09:42:41-0400

According to Newton's second law, the force is


"F=ma=m\cdot\frac{\Delta v}{\Delta t},"

where change in speed is


"\Delta v=v_2-v_1=v_0\text{cos}40^\circ -v_0=v_0(\text{cos}40^\circ-1)."

The flow of water "Q" during time "\Delta t" gives mass


"m=\rho Q\Delta t."

Substitution of the mass and the change in speed to the rightmost part of the first equation gives


"F=\rho Q\Delta t\cdot\frac{v_0(\text{cos}40^\circ-1)}{\Delta t}="

"=\rho Qv_0(\text{cos}40^\circ-1)=998\cdot0.28(0.766-1)=65.39\text{ N}."